93 71 What is the What is the 91 POWER What is the 16 18 What is the VAR S and ROUND

How many data observations are present (sample size)?

What is the sum of the data?

What is the mean (average) of the data (unrounded)?

What is the median of the data?

What is the minimum value of the data?

What is the maximum value of the data?

What is the square root of the sample size (unrounded)?

What is the value of the third observation when squared?

What is the square root of the sample size (to 2 digits)?

What is the square root of the sample size rounded down to the nearest whole number?

What is the square root of the sample size rounded up to the nearest whole number?

What is the sample variance (unrounded)?

What is the sample standard deviation (unrounded)?

What is the sample variance (to 2 decimals)?

What is the sample standard deviation (to 2 decimals)?

How many observations are less than 75?

Is the mean less than 75 (TRUE or FALSE)?

_______TRUEFALSE
Answer #1
count 11
sum 775
average 70.45455
median 65
min 52
max 93
sqrt 3.316625
power 4225
round 3.32

formulas

count =COUNT(A2:A12)
sum =SUM(A2:A12)
average =AVERAGE(A2:A12)
median =MEDIAN(A2:A12)
min =MIN(A2:A12)
max =MAX(A2:A12)
sqrt =SQRT(COUNT(A2:A12))
power =A4^2
round =ROUND(11,2)

Post rest parts again

> You didn't even finish it

mattgjg7 Thu, Feb 3, 2022 8:40 AM